# Stochastic heat flow by moments

### v1

@@ 4.1. Notation and tools
$\mathbb{R}^{4}\times\mathbb{R}^{2\omega\setminus\alpha}$ :=$:=\big{\{}y=(y_{\mathrm{r}},y_{\mathrm{c}},(y_{i})_{i\in\omega\setminus\alpha})\,\big{|}\,y_{c},y_{i}\in\mathbb{R}^{2}\big{\}}\ ,$ (4.2)
@@
Recall $\Phi$ from (1.3), set $\phi:=\sqrt{\Phi}$, and view $\Phi$ and $\phi$ as multiplicative operators acting on $L^{2}(\mathbb{R}^{2})=L^{2}(\{y_{\mathrm{r}}\})$. For $\alpha,\alpha^{\prime}\in\mathrm{Pair}(\omega)$, define operators
@@
$\begin{split}\mathcal{R}^{[\![n]\!],\varepsilon}_{\vec{\alpha}}(t):=\int_{\Sigma(t)}\mathrm{d}\vec{\tau}\,\mathcal{P}^{[\![n]\!],\varepsilon}_{\alpha_{1}}(\tau_{\frac{1}{2}})^{*}&\prod_{k=1}^{|\vec{\alpha}|-1}\Big{(}\beta_{\varepsilon}\delta_{0}(\tau_{k})+\mathcal{J}^{[\![n]\!],\varepsilon}_{\alpha_{k}}(\tau_{k})\Big{)}\,\mathcal{P}^{[\![n]\!],\varepsilon}_{\alpha_{k}\alpha_{k+1}}(\tau_{k+\frac{1}{2}})\\ &\cdot\Big{(}\beta_{\varepsilon}\delta_{0}(\tau_{|\vec{\alpha}|})+\mathcal{J}^{[\![n]\!],\varepsilon}_{\alpha_{|\vec{\alpha}|}}(\tau_{|\vec{\alpha}|})\Big{)}\,\mathcal{P}^{[\![n]\!],\varepsilon}_{\alpha_{|\vec{\alpha}|}}(\tau_{|\vec{\alpha}|+\frac{1}{2}})\ .\end{split}$ (4.6)
@@
We have slightly abused notation by using the same notation for the norms in (2.6) and in (4.7). This should not cause any confusion, since the latter will only be applied to $\varepsilon$-dependent operators. Define the operator norm $\|\cdot\|_{p,a\to p^{\prime},a^{\prime}}$ the same way as in (2.7) with the norm in (4.7) replacing that of (2.6). The following bounds are proven in [ST24, Section 3.1]. Given any finite $\omega\subset\mathbb{Z}_{>0}$, there exists $c=c(\phi,|\omega|,a)$ such that for all $t>0$ and $\alpha\neq\alpha^{\prime}\in\mathrm{Pair}(\omega)$,
@@
The solution of the mollified SHE enjoys a moment formula. Recall that $\mathsf{Z}^{\varepsilon}_{s,t}(x^{\prime},x)$ denotes the fundamental solution of (1.5), and let $\mathsf{W}^{\varepsilon}_{s,t}(x^{\prime},x):=\mathsf{Z}^{\varepsilon}_{s,t}(x^{\prime},x)-\mathsf{p}(t-s,x^{\prime}-x)$. For every $s<t$, $n\in\mathbb{Z}_{>0}$, $x=(x_{1},\ldots,x_{n}),x^{\prime}=(x^{\prime}_{1},\ldots,x^{\prime}_{n})\in\mathbb{R}^{2[\![n]\!]}$,
$\mathbf{E}\Big{[}\prod_{i=1}^{n}\mathsf{Z}^{\varepsilon}_{s,t}(x_{i},x^{\prime}_{i})\Big{]}=\mathcal{Q}^{[\![n]\!],\varepsilon}(t-s,x,x^{\prime})\ ,\ \mathbf{E}\Big{[}\prod_{i=1}^{n}\mathsf{W}^{\varepsilon}_{s,t}(x_{i},x^{\prime}_{i})\Big{]}=\mathcal{W}^{[\![n]\!],\varepsilon}(t-s,x,x^{\prime})\ .$ (4.10)
To prove (4.10), for $\alpha=ij$ and $x\in\mathbb{R}^{2[\![n]\!]}$, let $\Phi^{\varepsilon}_{\alpha}(x):=\beta_{\varepsilon}\Phi((x_{i}-x_{j})/\varepsilon)\varepsilon^{-2}$ and assume $s=0$ to simplify notation. By the arguments in [HN09, Theorem 5.3], moments of $\mathsf{Z}^{\varepsilon}_{s,t}(x,x^{\prime})$ enjoys the Feynman–Kac formula
$\mathbf{E}\Big{[}\prod_{i=1}^{n}\mathsf{Z}^{\varepsilon}_{0,t}(x_{i},x^{\prime}_{i})\Big{]}=\mathbb{E}_{\mathrm{BM}}\Big{[}\exp\Big{(}\int_{0}^{t}\mathrm{d}u\sum_{\eta\in\mathrm{Pair}[\![n]\!]}\Phi^{\varepsilon}_{\eta}(X(u))\Big{)}\prod_{i=1}^{n}\delta_{x^{\prime}_{i}}(X_{i}(t))\Big{]}\ ,$ (4.11)
where $X_{1},\ldots,X_{n}$ are independent Brownian motions on $\mathbb{R}^{2}$ with $X_{i}(s)=x_{i}$. Taylor expand the exponential in (4.11). The $\ell$th term of the result involves the integral $\frac{1}{\ell!}\prod_{i=1}^{\ell}\int_{s}^{t}\mathrm{d}u_{i}$. Rewrite the integral as $\int_{0<u_{1}<\ldots<u_{\ell}<t}\mathrm{d}\vec{u}$, rename $u_{i}-u_{i-1}$ to $s_{i}$, and further rewrite the integral as $\int_{\Sigma(t)}\mathrm{d}\vec{s}$. The result can be written as the kernel of
@@ 4.2. Checking the limiting properties
Let us check that every limit point of $\{Z^{\varepsilon}\}_{\varepsilon}$ satisfies Definition 1.2(b)(d). Part (d) follows from Proposition 1.6. Part (c) follows because $Z^{\varepsilon}_{s,t}$ and $Z^{\varepsilon}_{t,u}$ are independent, which holds because they are respectively measurable with respect to the sigma algebras generated by the noise $\xi$ on $[s,t]\times\mathbb{R}^{2}$ and on $[t,u]\times\mathbb{R}^{2}$. Finally, to check Part (b), fixing any $L^{2}$ applicable mollifier $\{u_{\ell}\}_{\ell}$, $2n\in\mathbb{Z}_{>0}$, and $g,g^{\prime}\in C_{\mathrm{c}}(\mathbb{R}^{2})$, we seek to prove that
$\lim_{\ell\to\infty}\lim_{\varepsilon\to 0}\mathbf{E}\big{[}\big{(}\big{(}Z^{\varepsilon}_{s,u}-Z^{\varepsilon}_{s,t}\bullet_{u_{\ell}}Z^{\varepsilon}_{t,u}\big{)}g\otimes g^{\prime}\big{)}^{2n}\big{]}=0\ .$ (4.13)
Once proven, this equality together with Proposition 1.6 implies that every limit point of $\{Z^{\varepsilon}\}_{\varepsilon}$ satisfies Part (b). To prove (4.13), first expand the expectation in (4.13) into
$\mathbf{E}\big{[}\big{(}\big{(}Z^{\varepsilon}_{s,u}-Z^{\varepsilon}_{s,t}\bullet_{u_{\ell}}Z^{\varepsilon}_{t,u}\big{)}(g\otimes g^{\prime})\big{)}^{2n}\big{]}=\sum_{k=0}^{2n}\binom{2n}{k}(-1)^{k}A_{k,\ell,\varepsilon}\ ,$ (4.14)
where $A_{k,\ell,\varepsilon}:=\mathbf{E}[(Z^{\varepsilon}_{s,u}g\otimes g^{\prime})^{2n-k}(Z^{\varepsilon}_{s,t}\bullet_{u_{\ell}}Z^{\varepsilon}_{t,u}g\otimes g^{\prime})^{2n-k}].$ Next, note that, since $\mathsf{Z}^{\varepsilon}_{s,t}(x,x^{\prime})$ is the fundamental solution of (1.5), for every $s<t<u$,
$Z^{\varepsilon}_{s,u}:=\mathrm{d}x\otimes\mathrm{d}x^{\prime}\,\mathsf{Z}^{\varepsilon}_{s,u}(x,x^{\prime})=\mathrm{d}x\otimes\mathrm{d}x^{\prime}\,\Big{(}\int_{\mathbb{R}^{2}}\mathrm{d}x^{\prime\prime}\,\mathsf{Z}^{\varepsilon}_{s,t}(x,x^{\prime\prime})\mathsf{Z}^{\varepsilon}_{t,u}(x^{\prime\prime},x^{\prime})\Big{)}=:Z^{\varepsilon}_{s,t}\,Z^{\varepsilon}_{t,u}\ .$ (4.15)
Use (4.15), (4.10), and the independence of $Z^{\varepsilon}_{s,t}$ and $Z^{\varepsilon}_{t,u}$ to evaluate $A_{k,\ell,\varepsilon}$. Doing so gives
@@
where $u^{\star}_{\ell}$ acts on $L^{2}(\mathbb{R}^{2})$ by convolution $u^{\star}_{\ell}\,\psi:=u_{\ell}\star\psi$. Using (4.9) for $(p,a)=(2,0)$ and Proposition 1.6 to take the $\varepsilon\to 0$ limit gives $\langle g^{\otimes 2n},\mathcal{Q}^{[\![2n]\!]}(t-s)\,(u_{\ell}^{\star}{}^{\otimes k}\otimes\mathbf{1}^{\otimes(2n-k)})\mathcal{Q}^{[\![2n]\!]}(u-t)\ ,g^{\prime}{}^{\otimes 2n}\rangle$. Since $\{u_{\ell}\}_{\ell}$ is an $L^{2}$ applicable mollifier, $u_{\ell}^{\star}\to\mathbf{1}$ strongly on $L^{2}(\mathbb{R}^{2})$ as $\ell\to\infty$. Hence
@@ 4.3. Tightness
For any $2n\in 2\mathbb{Z}_{>0}$, $p\in(2,\infty)$, and $a\in(0,\infty)$, with $11/p+1/q=$ 1,
$\big{\|}\mathcal{W}^{[\![2n]\!],\varepsilon}(t)\big{\|}_{p,a\to 2}=\big{\|}\mathcal{W}^{[\![2n]\!],\varepsilon}(t)\big{\|}_{2\to q,a}\leq c\,e^{ct}t^{\frac{n}{2}(1-\frac{2}{p})}\ ,\qquad c=c(n,\varphi,p,a)\ .$ (4.18)
@@
$\mathcal{B}^{\varepsilon}_{\alpha}(t):=\sum_{\vec{\eta}\in\mathrm{Dgm}_{\uparrow}[\![2n]\!],\eta_{n}=\alpha}\int_{\Sigma(t)}\mathrm{d}\vec{s}\,\mathcal{B}^{n,\varepsilon}_{\vec{\eta}}(s_{1},\ldots,s_{n})\otimes\mathcal{P}^{[\![2n]\!]\setminus(\eta_{1}\cup\cdots\cup\eta_{n})}(s_{1}+\ldots+s_{n})\ .$ (4.21)
The same argument that leads to (3.19) gives its $\varepsilon$ analog
$\mathcal{W}^{[\![2n]\!],\varepsilon}(t)=\sum_{\alpha\in\mathrm{Pair}[\![2n]\!]}\int_{s+s^{\prime}=t}\mathrm{d}s\,\,\mathcal{B}^{\varepsilon}_{\alpha}(s)\,\mathcal{C}^{[\![2n]\!],\varepsilon}_{\alpha}(s^{\prime})\ .$ (4.22)
@@
where $f_{1}(x^{\prime}):=(\int_{\mathbb{R}^{2}}\mathrm{d}x_{i_{*}}|f(x)|^{p})^{1/p}$, $x^{\prime}\in\mathbb{R}^{2\omega\setminus\{i_{*}\}}$, and $g_{1}(y):=\int_{\mathbb{R}^{2}}\mathrm{d}y_{\mathrm{r}}\,\phi(y_{\mathrm{r}})g(y_{\mathrm{r}},y_{\mathrm{c}}-\varepsilon y_{\mathrm{r}}/2,y_{\omega_{\ell}\setminus\alpha})$. Since $\|f_{1}\|_{p}=\|f\|_{p}\leq\|f\|_{p,a}$ (by definition) and since $\|g_{1}\|_{2}\leq\|g\|_{2}$ (by Cauchy–Schwarz), we have $\|\mathcal{B}^{\ell,\varepsilon}_{\vec{\alpha}}(\vec{s}\,)\|_{2\to q,a}\leq c\,(s_{1}+\ldots+s_{\ell})^{-1/p}\|\mathcal{B}^{\ell-1,\varepsilon}_{\vec{\eta}}(\vec{s})\,\mathcal{C}^{\omega_{\ell-1},\varepsilon}_{\eta_{\ell-1}}(s_{\ell})\|_{2\to q,a}$. Further using (4.23) with $(\omega,\alpha)=(\omega_{\ell-1},\eta_{\ell-1})$ gives
@@
The same bound holds in Case 2 by applying (4.24) with $(\omega,\alpha)=(\eta_{\ell},\eta_{\ell})$ and (4.23) with $(\omega,\alpha)=(\eta_{\ell-1},\eta_{\ell-1})$ to (4.20). In (4.29), bounding $(s_{1}+\ldots+s_{\ell})^{-1/p}\leq s_{\ell}^{-1/p}$ and using the induction hypothesis conclude (4.27).
@@
$\big{\|}\mathcal{B}^{\varepsilon}_{\alpha}(t)\big{\|}_{2\to q,a}\leq c\,e^{ct}\sum_{\vec{\eta}\in\mathrm{Dgm}_{\uparrow}[\![2n]\!],|\eta_{n}|=\alpha}\int_{s_{1}+\ldots+s_{n}=t}\mathrm{d}s_{1}\cdots\mathrm{d}s_{n-1}(s_{1}\cdots s_{n})^{-1/p}\,(s_{2}\cdots s_{n})^{-1/2}\ .$ (4.30)
Evaluate this integral and bound the size of the sum by $c(n)=c$ (see (3.13)) to get $\|\mathcal{B}^{\varepsilon}_{\alpha}(t)\|_{2\to q,a}\leq c\,e^{ct}t^{(n-1)/2-n/p}$. Insert this bound and (4.23) for $\omega=[\![n]\!]$ into (4.22), evaluate the integral, and bound the size of the sum by $c(n)=c$. Doing so gives the desired result (4.18). ∎
To prove the tightness of $\{Z^{\varepsilon}\}_{\varepsilon}$, recall that $D\subset C_{\mathrm{c}}^{2}(\mathbb{R}^{2})$ is countable and satisfies (2.21). Similar to Mitoma’s theorem, the following criterion holds, which we prove in the appendix. Equip $C(\mathbb{R}^{2}_{\scriptscriptstyle\leq},\mathbb{R})$ with the uniform-over-compact topology.
@@
(The law of) a set of $C(\mathbb{R}^{2}_{\scriptscriptstyle\leq},M_{+}(\mathbb{R}^{4}))$-valued random variables $\{Z^{\varepsilon}\}_{\varepsilon\in(0,1]}$ is tight if $\{Z^{\varepsilon}g\otimes g^{\prime}\}_{\varepsilon}$ is tight in $C(\mathbb{R}^{2}_{\scriptscriptstyle\leq},\mathbb{R})$ for every $g,g^{\prime}\in D$.
Given Lemma 4.2, it suffices to prove that, for every $g,g^{\prime}\in D$, $\{Z^{\varepsilon}g\otimes g^{\prime}\}_{\varepsilon}$ is tight in $C(\mathbb{R}^{2}_{\scriptscriptstyle\leq},\mathbb{R})$. This, by the multi-dimensional version of the Kolmogorov–Chentsov theorem (see [Kun90, Theorem 1.4.1] for example), follows from the moment estimate below.
@@
For every $2n\in 2\mathbb{Z}_{>0}$, $p\in(2,\infty)$, compact interval $I\subset\mathbb{R}$, and $g,g^{\prime}\in D$, there exists $c=c(n,\varphi,p,I,g,g^{\prime})$ such that, for all $[s,t],[s^{\prime},t^{\prime}]\subset I$,
$\big{(}\mathbf{E}\big{|}\big{(}Z^{\varepsilon}_{s,t}-Z^{\varepsilon}_{s^{\prime},t^{\prime}}\big{)}g\otimes g^{\prime}\big{|}^{2n}\big{)}^{1/2n}\leq c\,(|t-t^{\prime}|^{\frac{1}{2}(1-\frac{2}{p})}+|s-s^{\prime}|^{\frac{1}{2}(1-\frac{2}{p})})\ .$ (4.31)

### v3

@@ 4.1. Notation and tools
$\mathbb{R}^{4}\times\mathbb{R}^{2\omega\setminus\alpha}$ $:=\big\{y=(y_{\mathrm{r}},y_{\mathrm{c}},(y_{i})_{i\in\omega\setminus\alpha})\,\big|\,y_{\mathrm{r}},y_{\mathrm{c}},y_{i}\in\mathbb{R}^{2}\big\}\ ,$ (4.2)
@@
Recall $\Phi$ from (1.5), set $\phi:=\sqrt{\Phi}$, and view $\Phi$ and $\phi$ as multiplicative operators acting on $L^{2}(\mathbb{R}^{2})=L^{2}(\{y_{\mathrm{r}}\})$. For $\alpha,\alpha^{\prime}\in\mathrm{Pair}(\omega)$, define the operators
@@
$\begin{split}\mathcal{R}^{[\![n]\!],\varepsilon}_{\vec{\alpha}}(t):=\int_{\Sigma(t)}\hskip-5.0pt\mathrm{d}\vec{\tau}\,\mathcal{P}^{[\![n]\!],\varepsilon}_{\alpha_{1}}(\tau_{\frac{1}{2}})^{*}&\prod_{k=1}^{|\vec{\alpha}|-1}\Big(\beta_{\varepsilon}\delta_{0}(\tau_{k})+\mathcal{J}^{[\![n]\!],\varepsilon}_{\alpha_{k}}(\tau_{k})\Big)\,\mathcal{P}^{[\![n]\!],\varepsilon}_{\alpha_{k}\alpha_{k+1}}(\tau_{k+\frac{1}{2}})\\ &\cdot\Big(\beta_{\varepsilon}\delta_{0}(\tau_{|\vec{\alpha}|})+\mathcal{J}^{[\![n]\!],\varepsilon}_{\alpha_{|\vec{\alpha}|}}(\tau_{|\vec{\alpha}|})\Big)\,\mathcal{P}^{[\![n]\!],\varepsilon}_{\alpha_{|\vec{\alpha}|}}(\tau_{|\vec{\alpha}|+\frac{1}{2}})\ .\end{split}$ (4.6)
@@
We have slightly abused notation by using the same notation for the norms in (2.6) and in (4.7). This should not cause confusion, since the latter will only be applied to $\varepsilon$-dependent operators. Define the operator norm $\|\cdot\|_{p,a\to p^{\prime},a^{\prime}}$ the same way as in (2.7) with the norm in (4.7) replacing that of (2.6). The following bounds are proven in [ST25, Section 3.1]. Given any finite $\omega\subset\mathbb{Z}_{>0}$, there exists $c=c(\phi,|\omega|,a)$ such that for all $t>0$ and $\alpha\neq\alpha^{\prime}\in\mathrm{Pair}(\omega)$,
@@
The solution of the mollified SHE enjoys a moment formula. Recall that $\mathsf{Z}^{\theta,\varepsilon}_{s,t}(x^{\prime},x)$ denotes the fundamental solution of (1.7), and let $\mathsf{W}^{\varepsilon}_{s,t}(x^{\prime},x):=\mathsf{Z}^{\theta,\varepsilon}_{s,t}(x^{\prime},x)-\mathsf{p}(t-s,x^{\prime}-x)$. For every $s<t$, $n\in\mathbb{Z}_{>0}$, $x=(x_{1},\ldots,x_{n}),x^{\prime}=(x^{\prime}_{1},\ldots,x^{\prime}_{n})\in\mathbb{R}^{2[\![n]\!]}$,
$\mathbf{E}\Big[\prod_{i=1}^{n}\mathsf{Z}^{\theta,\varepsilon}_{s,t}(x_{i},x^{\prime}_{i})\Big]=\mathcal{Q}^{[\![n]\!],\varepsilon}(t-s,x,x^{\prime})\ ,\ \ \mathbf{E}\Big[\prod_{i=1}^{n}\mathsf{W}^{\varepsilon}_{s,t}(x_{i},x^{\prime}_{i})\Big]=\mathcal{W}^{[\![n]\!],\varepsilon}(t-s,x,x^{\prime})\ .$ (4.10)
To prove (4.10), for $\alpha=ij$ and $x\in\mathbb{R}^{2[\![n]\!]}$, let $\Phi^{\varepsilon}_{\alpha}(x):=\beta_{\varepsilon}\Phi((x_{i}-x_{j})/\varepsilon)\varepsilon^{-2}$ and assume $s=0$ to simplify notation. By the arguments in [HN09, Theorem 5.3], moments of $\mathsf{Z}^{\theta,\varepsilon}_{s,t}(x,x^{\prime})$ enjoys the Feynman–Kac formula
$\mathbf{E}\Big[\prod_{i=1}^{n}\mathsf{Z}^{\theta,\varepsilon}_{0,t}(x_{i},x^{\prime}_{i})\Big]=\mathbb{E}_{\mathrm{BM}}\Big[\exp\Big(\int_{0}^{t}\mathrm{d}u\sum_{\eta\in\mathrm{Pair}[\![n]\!]}\Phi^{\varepsilon}_{\eta}(X(u))\Big)\prod_{i=1}^{n}\delta_{x^{\prime}_{i}}(X_{i}(t))\Big]\ ,$ (4.11)
[…]
@@ 4.2. Checking the limiting properties
Let us check that every limit point of $\{Z^{\theta,\varepsilon}\}_{\varepsilon}$ satisfies Definition 1.2(b)(d). Part (d) follows from Proposition 1.7. Part (c) follows because $Z^{\theta,\varepsilon}_{t_{0},t_{1}},\ldots,Z^{\theta,\varepsilon}_{t_{k-1},t_{k}}$ are independent, which holds because they are respectively measurable with respect to the sigma algebras generated by the noise $\xi$ restricted to $t_{0},t_{1}]\times\mathbb{R}^{2},\ldots,[t_{k-1},t_{k}]\times\mathbb{R}^{2}$. Finally, to check Part (b), fixing any $L^{2}$ applicable mollifier $\{u_{\ell}\}_{\ell}$, $2n\in\mathbb{Z}_{>0}$, and $g,g^{\prime}\in C_{\mathrm{c}}(\mathbb{R}^{2})$, we seek to prove that
$\lim_{\ell\to\infty}\lim_{\varepsilon\to 0}\mathbf{E}\big[\big(\big(Z^{\theta,\varepsilon}_{s,u}-Z^{\theta,\varepsilon}_{s,t}\bullet_{u_{\ell}}Z^{\theta,\varepsilon}_{t,u}\big)g\otimes g^{\prime}\big)^{2n}\big]=0\ .$ (4.13)
Once proven, this equality together with Proposition 1.7 implies that every limit point of $\{Z^{\theta,\varepsilon}\}_{\varepsilon}$ satisfies Part (b). To prove (4.13), first expand the expectation in (4.13) into
$\mathbf{E}\big[\big(\big(Z^{\theta,\varepsilon}_{s,u}-Z^{\theta,\varepsilon}_{s,t}\bullet_{u_{\ell}}Z^{\theta,\varepsilon}_{t,u}\big)(g\otimes g^{\prime})\big)^{2n}\big]=\sum_{k=0}^{2n}\binom{2n}{k}(-1)^{k}A_{k,\ell,\varepsilon}\ ,$ (4.14)
where $A_{k,\ell,\varepsilon}:=\mathbf{E}[(Z^{\theta,\varepsilon}_{s,u}g\otimes g^{\prime})^{2n-k}(Z^{\theta,\varepsilon}_{s,t}\bullet_{u_{\ell}}Z^{\theta,\varepsilon}_{t,u}g\otimes g^{\prime})^{2n-k}].$ Next, note that, since $\mathsf{Z}^{\theta,\varepsilon}_{s,t}(x,x^{\prime})$ is the fundamental solution of (1.7), for every $s<t<u$,
$Z^{\theta,\varepsilon}_{s,u}:=\mathrm{d}x\otimes\mathrm{d}x^{\prime}\,\mathsf{Z}^{\theta,\varepsilon}_{s,u}(x,x^{\prime})=\mathrm{d}x\otimes\mathrm{d}x^{\prime}\,\Big(\int_{\mathbb{R}^{2}}\mathrm{d}x^{\prime\prime}\,\mathsf{Z}^{\theta,\varepsilon}_{s,t}(x,x^{\prime\prime})\mathsf{Z}^{\theta,\varepsilon}_{t,u}(x^{\prime\prime},x^{\prime})\Big)=:Z^{\theta,\varepsilon}_{s,t}\,Z^{\theta,\varepsilon}_{t,u}\ .$ (4.15)
Use (4.15), (4.10), and the independence of $Z^{\theta,\varepsilon}_{s,t}$ and $Z^{\theta,\varepsilon}_{t,u}$ to evaluate $A_{k,\ell,\varepsilon}$. Doing so gives
@@
where $u^{\star}_{\ell}$ acts on $L^{2}(\mathbb{R}^{2})$ by convolution $u^{\star}_{\ell}\,\psi:=u_{\ell}\star\psi$. Using (4.9) and Proposition 1.7 to take the $\varepsilon\to 0$ limit gives $\langle g^{\otimes 2n},\mathcal{Q}^{[\![2n]\!]}(t-s)\,(u_{\ell}^{\star}{}^{\otimes k}\otimes\mathbf{1}^{\otimes(2n-k)})\mathcal{Q}^{[\![2n]\!]}(u-t)\ ,g^{\prime}{}^{\otimes 2n}\rangle$. Since $\{u_{\ell}\}_{\ell}$ is an $L^{2}$ applicable mollifier, $u_{\ell}^{\star}\to\mathbf{1}$ strongly on $L^{2}(\mathbb{R}^{2})$ as $\ell\to\infty$. Hence
@@ 4.3. Tightness
For any $2n\in 2\mathbb{Z}_{>0}$, $p\in(2,\infty)$, and $a\in(0,\infty)$, with $1/p+1/q=1$,
$\big\|\mathcal{W}^{[\![2n]\!],\varepsilon}(t)\big\|_{p,a\to 2}=\big\|\mathcal{W}^{[\![2n]\!],\varepsilon}(t)\big\|_{2\to q,a}\leq c\,e^{ct}t^{\frac{n}{2}(1-\frac{2}{p})}\ ,\qquad c=c(n,\phi,p,a)\ .$ (4.18)
@@
$\widetilde{\mathcal{B}}^{\varepsilon}_{\vec{\eta}}(t)$ $:=\int_{\Sigma(t)}\mathrm{d}\vec{s}\,\mathcal{B}^{n,\varepsilon}_{\vec{\eta}}(s_{1},\ldots,s_{n})\otimes\mathcal{P}^{[\![2n]\!]\setminus(\eta_{1}\cup\cdots\cup\eta_{n})}(s_{1}+\ldots+s_{n})\ ,$ (4.21)
$\widetilde{\mathcal{C}}^{\varepsilon}_{\vec{\eta}}(t)$ $:=\sum\int_{\Sigma(s)}\mathrm{d}\vec{\tau}\,\prod_{k=1}^{|\vec{\gamma}|-1}\mathcal{J}^{\omega,\varepsilon}_{\gamma_{k}}(\tau_{k})\,\mathcal{P}^{\omega,\varepsilon}_{\gamma_{k}\gamma_{k+1}}(\tau_{k+1/2})\,\cdot\mathcal{J}^{\omega,\varepsilon}_{\gamma_{|\vec{\gamma}|}}(\tau_{|\vec{\gamma}|})\,\mathcal{P}^{\omega,\varepsilon}_{\gamma_{|\vec{\gamma}|}}(\tau_{|\vec{\gamma}|+1/2})\ ,$ (4.22)
where the last sum runs over $\vec{\gamma}\in\mathrm{Dgm}[\![n]\!]$ under the constraint that $\eta_{1}\cup\cdots\eta_{n}\cup\gamma_{1}\cup\cdots\gamma_{|\gamma|}=[\![2n]\!]$. The same argument that leads to (3.20) gives its $\varepsilon$ analog
$\mathcal{W}^{[\![2n]\!],\varepsilon}(t)=\sum_{\vec{\eta}\in\mathrm{Dgm}_{\uparrow}[\![2n]\!]}\int_{s+s^{\prime}=t}\mathrm{d}s\,\,\widetilde{\mathcal{B}}^{\varepsilon}_{\vec{\eta}}(s)\,\widetilde{\mathcal{C}}^{\varepsilon}_{\vec{\eta}}(s^{\prime})\ .$ (4.23)
@@
where $f_{1}(x^{\prime}):=(\int_{\mathbb{R}^{2}}\mathrm{d}x_{i_{*}}|f(x)|^{p})^{1/p}$, $x^{\prime}\in\mathbb{R}^{2\omega\setminus\{i_{*}\}}$, and $g_{1}(y):=\int_{\mathbb{R}^{2}}\mathrm{d}y_{\mathrm{r}}\,\phi(y_{\mathrm{r}})g(y_{\mathrm{r}},y_{\mathrm{c}}-\varepsilon y_{\mathrm{r}}/2,y_{\omega_{\ell}\setminus\alpha})$. Since $\|f_{1}\|_{p}=\|f\|_{p}\leq\|f\|_{p,a}$ (by definition) and since $\|g_{1}\|_{2}\leq\|\phi\|_{2}\|g\|_{2}=\|g\|_{2}$ (by Cauchy–Schwarz), we have $\|\mathcal{B}^{\ell,\varepsilon}_{\vec{\alpha}}(\vec{s}\,)\|_{2\to q,a}\leq c\,(s_{1}+\ldots+s_{\ell})^{-1/p}\|\mathcal{B}^{\ell-1,\varepsilon}_{\vec{\eta}}(\vec{s})\,\mathcal{C}^{\omega_{\ell-1},\varepsilon}_{\eta_{\ell-1}}(s_{\ell})\|_{2\to q,a}$. Further using (4.24) with $(\omega,\alpha)=(\omega_{\ell-1},\eta_{\ell-1})$ gives
@@
The same bound holds in Case 2 by applying (4.25) with $(\omega,\alpha)=(\eta_{\ell},\eta_{\ell})$ and (4.24) with $(\omega,\alpha)=(\omega_{\ell-1},\eta_{\ell-1})$ to (4.20). In (4.30), bounding $(s_{1}+\ldots+s_{\ell})^{-1/p}\leq s_{\ell}^{-1/p}$ and using the induction hypothesis conclude (4.28).
@@
$\big\|\widetilde{\mathcal{B}}^{\varepsilon}_{\alpha}(t)\big\|_{2\to q,a}\leq c\,e^{ct}\int_{s_{1}+\ldots+s_{n}=t}\hskip-20.0pt\mathrm{d}s_{1}\cdots\mathrm{d}s_{n-1}(s_{1}\cdots s_{n})^{-1/p}\,(s_{2}\cdots s_{n})^{-1/2}\ .$ (4.31)
Evaluate this integral gives the bound $\leq c\,e^{ct}t^{(n-1)/2-n/p}$. Insert this bound and (4.24) into (4.23), evaluate the integral, and bound the size of the sum y $c(n)=c$ (see (3.13)). Doing so gives the desired result (4.18). ∎
To prove the tightness of $\{Z^{\theta,\varepsilon}\}_{\varepsilon}$, recall that $D\subset C_{\mathrm{c}}^{2}(\mathbb{R}^{2})$ is countable and satisfies (2.21). Similar to Mitoma’s theorem, the following criterion holds, which we prove in the appendix. Equip $C(\mathbb{R}^{2}_{\scriptscriptstyle\leq},\mathbb{R})$ with the uniform-over-compact topology.
@@
(The law of) a set of $C(\mathbb{R}^{2}_{\scriptscriptstyle\leq},\mathcal{M}_{+}(\mathbb{R}^{4}))$-valued random variables $\{Z^{\theta,\varepsilon}\}_{\varepsilon\in(0,1]}$ is tight if $\{Z^{\theta,\varepsilon}g\otimes g^{\prime}\}_{\varepsilon}$ is tight in $C(\mathbb{R}^{2}_{\scriptscriptstyle\leq},\mathbb{R})$ for every $g,g^{\prime}\in D$.
Given Lemma 4.2, it suffices to prove that, for every $g,g^{\prime}\in D$, $\{Z^{\theta,\varepsilon}g\otimes g^{\prime}\}_{\varepsilon}$ is tight in $C(\mathbb{R}^{2}_{\scriptscriptstyle\leq},\mathbb{R})$. This, by the multi-dimensional version of the Kolmogorov–Chentsov theorem (see [Kun90, Theorem 1.4.1] for example), follows from the moment estimate below.
@@
For every $2n\in 2\mathbb{Z}_{>0}$, $p\in(2,\infty)$, compact interval $I\subset\mathbb{R}$, and $g,g^{\prime}\in D$, there exists $c=c(n,\phi,p,I,g,g^{\prime})$ such that, for all $s,t],[s^{\prime},t^{\prime}]\subset I$,
$\big(\mathbf{E}\big|\big(Z^{\theta,\varepsilon}_{s,t}-Z^{\theta,\varepsilon}_{s^{\prime},t^{\prime}}\big)g\otimes g^{\prime}\big|^{2n}\big)^{1/2n}\leq c\,(|t-t^{\prime}|^{\frac{1}{4}(1-\frac{2}{p})}+|s-s^{\prime}|^{\frac{1}{4}(1-\frac{2}{p})})\ .$ (4.32)